OpenStudy (anonymous):

The ratio of copper and iron in an alloy is 4:5 .If there is 25% impurity in copper then find the amount of impurity in 18 kg of the alloy.

OpenStudy (kira_yamato):

Amount of Cu in the alloy = 18*4/9 = 8kg Since 1/4 of the Cu is impurity, Amount of impurity in alloy = 8*1/4 = 2kg